# Toward a Physical Theory of Intelligence

### v1

@@ 2.1 Agent-Environment Dynamics Model
Although the presentation in this subsection uses discrete-time stochastic kernels for clarity, the framework admits a fully continuous-time formulation. In particular, the one-step kernels $K_{X}$ and $K_{E}$ may be replaced by a joint Markov generator $\mathcal{L}$ acting on bounded measurable functions over $X\times E$, or equivalently by coupled stochastic differential equations whose drift and diffusion coefficients depend on the port variables $\Gamma_{X}(X_{t})$ and $\Gamma_{E}(E_{t})$. The resulting semigroup $e^{t\mathcal{L}}$ defines a family of transition kernels that recovers the discrete-time model as a special case. The GENERIC decomposition used later in the paper naturally operates in this continuous-time setting.
2.2 Goal Dynamics and Energy Potentials
Goals of the coupled agent-environment system are modelled by scalar potential functions admitting energetic interpretation,
$G:X\times E\to\mathbb{R},$ (9)
where $G(x,e)$ denotes the instantaneous potential energy of the joint state $S_{t}:=(X_{t},E_{t})$. For clarity, we initially restrict attention to conservative potentials, so that the work performed along a trajectory depends only on its endpoints. Under this assumption, and assuming sufficiently regular dynamics for which the path derivative $\dot{S}_{t}$ is well defined, the total goal-directed work over a finite horizon $T$ is
$W_{T}=-\,\mathbb{E}\!\left[\int_{0}^{T}\nabla_{(x,e)}G(S_{t})\cdot\dot{S}_{t}\,dt\right]=\mathbb{E}[G(S_{0})]-\mathbb{E}[G(S_{T})],$ (10)
where the equality follows from the fundamental theorem for line integrals applied to conservative fields.
A positive value of $W_{T}$ indicates conversion of stored potential energy into goal-directed work, while a negative value corresponds to divergence from the goal or accumulation of energetic debt. The associated goal functional is defined by
$\Psi(G\mid\mu_{t}):=-\!\int G(x,e)\,d\mu_{t}(x,e),$ (11)
where $\mu_{t}=\mathrm{Law}(S_{t})$ denotes the joint state distribution at time $t$. Restricting goals to energy-based potentials ensures that all subsequent quantities, including information and intelligence metrics, maintain consistent physical units and interpretation.
More general goal structures may be represented using non-conservative or explicitly time-dependent potentials $G(x,e,t)$, which capture evolving task landscapes, adaptive objectives, or externally imposed performance criteria. In such settings, goal-directed work need not correspond to the decrease of a conserved potential, but may instead be defined operationally via monotone task-aligned functionals whose variation tracks the system’s causal influence on task-relevant degrees of freedom. Provided these functionals are consistently coupled to the agent–environment dynamics, they induce a well-defined notion of useful work as the physically mediated reduction of task error, uncertainty, or constraint violation over time. This generalized formulation preserves the core interpretation of intelligence as the efficiency with which irreversible information processing is converted into goal-aligned influence, while allowing empirical instantiations in settings where an explicit conservative energy potential is unavailable or inappropriate.
2.3 Information, Encoding, and Conserved Quantities
Intelligent systems process information through physical substrates. To relate abstract notions of information and computation to measurable physical quantities, we formalize how encodings of information are instantiated in physical systems and how these encodings interact with conserved quantities. Encodings are represented by physical configurations that satisfy a general Conservation-Congruent Encoding (CCE) condition defined below. CCE guarantees that changes in entropy from an information-theoretic perspective correspond to changes in the entropy of an underlying physical measure constrained by a conservation law. This establishes the essential link between irreversible information processing and the expenditure of conserved resources such as energy, angular momentum, or chemical potential [18, 6, 21].
2.3.1 Information Encoding and Physical State
We distinguish between a set of encodings $L$ and a physical state space $P$. Let $C$ denote a collection of externally adjustable control parameters (e.g., barrier heights, biases, or couplings). A physical system realizes an encoding $\ell\in L$ through a probability distribution over physical states under a given control configuration $c\in C$. Formally, the encoding map
$f:L\times C\to\mathcal{P}(P)$ (12)
assigns to each pair $(\ell,c)$ a measure $f(\ell,c)$ over $P$, representing the set of physical configurations consistent with encoding $\ell$. To ensure that encodings are physically meaningful, we impose the Conservation-Congruent Encoding (CCE) conditions:
@@ 2.3.1 Information Encoding and Physical State
    Metastable separation: Each encoding $\ell\in L$ occupies a disjoint metastable basin $B_{\ell}\subset P$, separated from all others by an activation barrier associated with a conserved physical quantity. Within each basin, local equilibration occurs on timescales short compared to those on which basin structure or inter-basin transitions are modified [16, 15].
@@
    Quasistatic controllability: Control trajectories $c_{t}$ deform the basin structure $\{B_{\ell}\}$ quasistatically relative to internal relaxation. Transitions between basins proceed through well-defined paths whose statistics are constrained by the conserved-quantity structure of the underlying dynamics (e.g. via local detailed balance), rather than by arbitrary control-induced jumps [8, 24].
Under CCE, the encoding map $f$ is locally invertible within each metastable basin: physical configurations identify encodings up to fluctuations of order $O(\alpha)$, where $\alpha$ is the scale of entropy fluctuations associated with the relevant conserved quantity (e.g., $k_{B}$ for thermal channels). Consequently, variations in logical entropy are congruent with variations in the entropy of the associated physical measure,
$\Delta S_{\mathrm{phys}}\approx\alpha\,\Delta H_{L},$ (13)
where $H_{L}$ denotes the Shannon entropy of the encoding distribution. This congruence between logical entropy and physical entropy under irreversible operations generalizes Landauer’s principle beyond thermal reservoirs [18, 6, 21].
If CCE is violated, for instance, if several encodings collapse into the same metastable basin or if reversible dynamics fail to preserve the conserved quantity, then logical irreversibility may occur without yielding a detectable physical signature. Such encodings fall outside the scope of the present framework. Unless otherwise stated, all subsequent quantities assume CCE-compliant encodings. A formal example illustrating these assumptions is provided in Appendix A.

### v2

@@ 2.1 Agent–Environment Dynamics Model
2.2 Useful Work as a Spatiotemporal Causal Footprint
Intelligent agents typically operate in open, spatiotemporally extended environments. Small, targeted interventions at the agent–environment interface can unlock large downstream changes in the environment (e.g. flipping a single bit that triggers a large automated control action, initiating a reaction by introducing a catalyst, or issuing a command that coordinates distributed actuators). Evaluating intelligence therefore requires a macroscopic notion of causal footprint: the total downstream structural consequences causally enabled by the agent’s interventions.
Because explicitly simulating the environment $E$ at all microscopic scales is intractable, we evaluate downstream dynamics through a spatiotemporal initial–boundary value problem (IBVP) constrained by relativistic causality and local conservation laws.
Let $\mathcal{M}$ denote a spacetime manifold foliated by spatial hypersurfaces $\Sigma_{t}$. We formalise the agent’s spatial boundary as a closed, codimension-one interface $\Gamma_{\mathrm{port}}\subset\Sigma_{t}$. Over the horizon $t\in[0,T]$, this interface sweeps out a timelike world-tube
$\mathcal{W}_{\mathrm{port}}:=\Gamma_{\mathrm{port}}\times[0,T]\subset\mathcal{M}.$ (9)
This provides the physical bridge to the stochastic port model of §2.1: the abstract port variables $\Pi_{X}(X_{t})$ and $\Pi_{E}(E_{t})$ manifest physically as time-dependent boundary data exchanged across $\mathcal{W}_{\mathrm{port}}$.
We designate causal past and future using standard notation. The agent’s interventions at the port can only influence the environment within its future domain of dependence $J^{+}(\mathcal{W}_{\mathrm{port}})$. The historical data determining the initial state of the environment is strictly contained within the causal past $J^{-}(\mathcal{W}_{\mathrm{port}})$. At time $t\geq 0$, the downstream region of influence is the spatial volume
$\Omega(t):=\Sigma_{t}\cap J^{+}(\mathcal{W}_{\mathrm{port}}).$ (10)
Its boundary $\partial\Omega(t)$ is partitioned into two disjoint surfaces: (i) the active port interface $\Gamma_{\mathrm{port}}$, and (ii) the expanding causal horizon $\Gamma_{\mathrm{LC}}(t)$ separating the influenced region from the rest of the universe.
To remain substrate-neutral, we avoid committing to a single reference state. Instead, we model the continuum within $\Omega(t)$ via a generalised structural potential density $a(\mathbf{r},t)$ and its spatial flux $\mathbf{J}_{a}(\mathbf{r},t)$. To ensure dimensional consistency with the mesoscopic informational costs developed in subsequent sections, we formally parameterise this potential across the relevant physical channels as
$a(\mathbf{r},t)=\sum_{c}(\mathcal{F}_{c}\alpha_{c})\rho_{c}(\mathbf{r},t),$ (11)
where $\rho_{c}(\mathbf{r},t)$ is the local density of accessible structural distinctions (a dimensionless volumetric density) for channel $c$. The parameter $\alpha_{c}$ is the characteristic physical scale of that channel (e.g., $k_{B}$ for thermal limits), and $\mathcal{F}_{c}$ is the corresponding intensive conjugate force.
This multiplexed formulation preserves complete substrate independence. It defines $a(\mathbf{r},t)$ as the locally available structural capacity (a generalised free energy) evaluated at the chosen operational scale, while ensuring that the macroscopic causal footprint remains physically commensurate with the agent’s internal computational costs. Let $\sigma_{a}(\mathbf{r},t)\geq 0$ denote the corresponding local rate density of irreversible structural dissipation across the integrated channels.
To isolate the agent’s causal influence, the environmental evolution is posed as an IBVP with two constraints:
  1. (a)
    Initial Cauchy data: the state on $\Omega(0)$ is determined by the universe history in $J^{-}(\mathcal{W}_{\mathrm{port}})\cap\{t\leq 0\}$, so inherited pre-existing dynamics are explicitly conditioned upon.
  2. (b)
    Port boundary condition: along $\mathcal{W}_{\mathrm{port}}$, the agent prescribes the active fluxes injected into the environment.
We define the absolute enabled work over horizon $T$ as the integral of all downstream generalised dissipation within the causal region,
$W_{\mathrm{enabled},T}$ $:=\int_{0}^{T}dt\int_{\Omega(t)}\sigma_{a}(\mathbf{r},t)\,d^{3}\mathbf{r}.$ (12)
Local conservation laws and irreversible degradation imply a continuity balance
$\partial_{t}a+\nabla\cdot\mathbf{J}_{a}=-\sigma_{a}.$ (13)
Applying the Reynolds transport theorem and Gauss’s divergence theorem to the expanding volume $\Omega(t)$ relates volumetric structural decay to boundary fluxes:
$\frac{d}{dt}\int_{\Omega(t)}a(\mathbf{r},t)\,d^{3}\mathbf{r}$ $=\int_{\Gamma_{\mathrm{port}}}\mathbf{J}_{a}\cdot\mathbf{n}_{\mathrm{in}}\,dA-\int_{\Gamma_{\mathrm{LC}}(t)}\!\!\left(\mathbf{J}_{a}-a\,\mathbf{v}_{\mathrm{LC}}\right)\cdot\mathbf{n}_{\mathrm{out}}\,dA-\int_{\Omega(t)}\sigma_{a}(\mathbf{r},t)\,d^{3}\mathbf{r},$ (14)
where $\mathbf{n}_{\mathrm{in}}$ is the inward normal from the port and $\mathbf{v}_{\mathrm{LC}}$ is the expansion velocity of the causal horizon. Reynolds transport gives
$\frac{d}{dt}\int_{\Omega(t)}a\,d^{3}\mathbf{r}=\int_{\Omega(t)}\partial_{t}a\,d^{3}\mathbf{r}+\int_{\partial\Omega(t)}a\,\mathbf{v}_{\partial\Omega}\cdot\mathbf{n}_{\mathrm{out}}\,dA,$ (15)
with $\mathbf{v}_{\partial\Omega}=\mathbf{0}$ on the fixed port boundary and $\mathbf{v}_{\partial\Omega}=\mathbf{v}_{\mathrm{LC}}$ on the moving boundary $\Gamma_{\mathrm{LC}}(t)$. Substituting the local balance $\partial_{t}a=-\nabla\cdot\mathbf{J}_{a}-\sigma_{a}$ and applying Gauss’s theorem yields
$\frac{d}{dt}\int_{\Omega(t)}a\,d^{3}\mathbf{r}$ $=-\int_{\partial\Omega(t)}\mathbf{J}_{a}\cdot\mathbf{n}_{\mathrm{out}}\,dA+\int_{\partial\Omega(t)}a\,\mathbf{v}_{\partial\Omega}\cdot\mathbf{n}_{\mathrm{out}}\,dA-\int_{\Omega(t)}\sigma_{a}\,d^{3}\mathbf{r},$ (16)
which, after splitting $\partial\Omega(t)=\Gamma_{\mathrm{port}}\cup\Gamma_{\mathrm{LC}}(t)$ and using $\mathbf{n}_{\mathrm{in}}=-\mathbf{n}_{\mathrm{out}}$ on $\Gamma_{\mathrm{port}}$, gives the stated form. The first boundary term evaluates to the physical power injected by the agent, $\dot{W}_{\mathrm{port}}(t)$. Let $\Phi_{\mathrm{LC}}(t)$ denote the net flux escaping across the expanding causal horizon. Under causal isolation of the region, there is no incoming flux across the horizon, so the horizon term reduces to outward leakage only (and in the idealised perfectly isolating limit, $\Phi_{\mathrm{LC}}(t)=0$). Integrating over $[0,T]$ yields the exact relation
$W_{\mathrm{enabled},T}$ $=W_{\mathrm{port},T}-\Delta A_{\mathrm{env}}-\int_{0}^{T}\Phi_{\mathrm{LC}}(t)\,dt,$ (17)
where $W_{\mathrm{port},T}:=\int_{0}^{T}\dot{W}_{\mathrm{port}}(t)\,dt$ and $\Delta A_{\mathrm{env}}:=A_{\mathrm{env}}(T)-A_{\mathrm{env}}(0)$ is the net change in total bulk structural potential within $\Omega(t)$.
The initial data on $\Omega(0)$ already encode gradients inherited from the past light cone. Autonomous relaxation of these pre-existing gradients will produce structural change and flux across $\Gamma_{\mathrm{LC}}(t)$ even if the agent behaves as a strictly passive object. To avoid conflating active causal influence with deterministic background drift, define a baseline using the same past-light-cone conditioned initial data but imposing a passive boundary condition at the port world-tube (e.g. $\mathbf{J}_{a}\cdot\mathbf{n}=0$ on $\mathcal{W}_{\mathrm{port}}$). Let $W_{\mathrm{enabled},T}^{\mathrm{default}}$ and $\Delta A_{\mathrm{env}}^{\mathrm{default}}$ denote the enabled work and bulk structural change under this passive baseline.
By domain-of-dependence, the state on and outside the expanding causal horizon $\Gamma_{\mathrm{LC}}(t)$ cannot depend on boundary conditions applied at $\mathcal{W}_{\mathrm{port}}$ during $t\in[0,T]$. Therefore, the horizon flux $\Phi_{\mathrm{LC}}(t)$ is identical in the actual and default dynamics. Defining the purely causally enabled structural footprint as
$W_{\mathrm{causal},T}$ $:=W_{\mathrm{enabled},T}^{\mathrm{actual}}-W_{\mathrm{enabled},T}^{\mathrm{default}},$ (18)
the unwieldy light-cone terms cancel exactly:
$W_{\mathrm{causal},T}$ $=W_{\mathrm{port},T}-\Big(\Delta A_{\mathrm{env}}^{\mathrm{actual}}-\Delta A_{\mathrm{env}}^{\mathrm{default}}\Big)=W_{\mathrm{port},T}-\Delta A_{\mathrm{env}}^{\mathrm{causal}}.$ (19)
This boundary-value formulation isolates the agent’s net structural footprint without assuming an “unperturbed” external vacuum, and ensures an agent receives no efficiency credit for merely witnessing pre-existing cascades (a large autonomous background event yields $\Delta A_{\mathrm{env}}^{\mathrm{causal}}=0$).
Often we are interested not in total causal footprint but in influence aligned with a task or purpose at the chosen operational scale. To account for this a projection operator $\pi_{\mathcal{G}}$ that extracts the goal-relevant component of the interface-level footprint can be leveraged. We define the expected work over horizon $T$ as
$W_{\mathrm{goal},T}$ $:=\mathbb{E}_{\mu_{0:T}}\!\left[\pi_{\mathcal{G}}\!\left(W_{\mathrm{port},T}-\Delta A_{\mathrm{env}}^{\mathrm{causal}}\right)\right],$ (20)
where $\mu_{0:T}$ is the law of the coupled process on $[0,T]$.
2.3 Information Processing
Intelligent systems process information through the evolution of physical dynamics. Throughout, we adopt the language of information processing as a descriptive convenience. The underlying physical primitive is entropy production, which can be related (at an appropriate level of description) to information-theoretic quantities, but is not identical to them. To relate abstract notions of information to measurable quantities, we formalise information processing at the level of probability distributions over system trajectories. In this framework, information corresponds to dynamically preserved distinctions in the system’s evolution, while irreversible information loss corresponds to the collapse of such distinctions under time evolution.
In practice, irreversible information processing may be diagnosed via divergence between forward and time-reversed trajectory distributions. Such divergences provide bounds on irreversible loss of information, becoming informative only after projection onto a dynamically induced quotient space that determines which distinctions are physically meaningful.
While this path-space formulation applies to general dynamical systems, relating informational irreversibility to explicitly measurable physical costs typically requires additional structure. In particular, not every operational coarse-graining admits a clean metastable encoding at a given scale. We therefore introduce the Conservation-Congruent Encoding (CCE) framework, which specifies a class of systems in which dynamically preserved distinctions admit metastable physical realizations constrained by conservation laws. In this regime, information-theoretic measures of irreversibility can be directly related to entropy production and the dissipation of conserved quantities such as heat or angular momentum, recovering and generalising Landauer-type bounds [15, 5, 17].
2.3.1 Path-Space Information Measures for Dynamical Systems
Information processing arises from the irreversible evolution of coupled dynamical systems. In this work, we take the base state spaces $(X\times E,\mathscr{X}\otimes\mathscr{E})$ to represent the absolute microscopic reality of the agent and environment, governed fundamentally by deterministic, reversible dynamics. Consequently, the stochasticity present in the induced agent transition kernel $K_{X}(dx^{\prime}|x)$ is not an axiomatic assumption, but a strict mathematical artifact of the port maps $\Pi_{X}$ and $\Pi_{E}$. Because the agent interacts with the environment exclusively through these lower-dimensional boundaries, the unresolved microscopic degrees of freedom of the environment act as a bath, driving the irreversible stochastic evolution of the open subsystem $X$.
While instantaneous distributions $\mu_{t}^{X}=\mathrm{Law}(X_{t})$ describe the subsystem at a given time, questions of reversibility and irreversible information loss are inherently temporal and are therefore most naturally formulated at the level of entire trajectories. Let
$\mathcal{T}_{T}:=\{\tau=(x_{0},x_{1},...,x_{T})|x_{t}\in X\}$ (21)
denote the space of agent trajectories over a finite horizon $T$. Given an initial distribution $\mu_{0}^{X}\in\mathcal{P}(X)$ and the induced stochastic transition kernel $K_{X}$, the dynamics generate a probability measure over trajectories, $\mathbb{P}_{\mu_{0}^{X}}$, defined on $\mathcal{T}_{T}$.
Operational notions of information and computation may not depend on all microscopic distinctions between internal states. Furthermore, in complex systems, information processing is often multiplexed across multiple spatiotemporal scales and macroscopic physical channels. We therefore define a multiplexed operational interface via a projection $\Phi:=\{\phi_{t}:X\rightarrow Z\}_{t=0}^{T}$ where the target $Z$ acts as a quotient space grouping microscopic states into macroscopic equivalence classes. To align with the macroscopic properties of the environment, the operational space is strictly a Cartesian product of distinct physical channels, $Z=Z_{1}\times Z_{2}\times\dots\times Z_{C}$. Each component projection $\phi^{(c)}_{t}:X\rightarrow Z_{c}$ isolates the dynamical equivalence classes associated with a specific macroscopic conserved quantity (e.g., thermal heat, chemical gradients, or angular momentum).
This projection induces a corresponding trajectory-level mapping $\pi_{\Phi}:\mathcal{T}_{T}\rightarrow\mathcal{Z}_{T}$. The induced statistics of coarse-grained trajectories are captured by the pushforward measure $\mathbb{P}_{\mu_{0}^{X}}^{\Phi}:=(\pi_{\Phi})_{\#}\mathbb{P}_{\mu_{0}^{X}}$. To assess temporal asymmetry at this operational level, we introduce a time-reversal operator $\Theta$ and define the time-reversed induced path measure $(\mathbb{P}_{\mu_{0}^{X}}^{rev})^{\Phi}$.
We define the operational (or encoding-level) irreversibility over horizon $T$ as the statistical distinguishability between the forward and time-reversed induced path measures:
$\mathcal{I}_{\Phi}(T):=D_{KL}(\mathbb{P}_{\mu_{0}^{X}}^{\Phi}\,\|\,(\mathbb{P}_{\mu_{0}^{X}}^{rev})^{\Phi}).$ (22)
Operational reversibility corresponds to the vanishing of this quantity ($\mathcal{I}_{\Phi}(T)=0$). This formulation ensures invariance under internal permutations, reparameterizations, or spatial translations within the system. For example, if an encoded predictive memory is physically routed from one spatial region of the agent to another, the projection identifies these distinct micro-trajectories as belonging to the same preserved equivalence class, correctly yielding zero irreversible computational cost for pure spatial transport.
By the data-processing inequality, encoding-level irreversibility is bounded above by the raw trajectory-level asymmetry of the open system:
$\mathcal{I}_{\Phi}(T)\leq\mathcal{I}(T).$ (23)
Crucially, because $X$ is an open system coupled to an unobserved environment via the port maps, its induced microscopic evolution is non-unitary. Therefore, the base trajectory asymmetry $\mathcal{I}(T)$ is strictly positive, representing the total physical entropy exported to the environmental bath. The data-processing inequality thus asserts a strict physical hierarchy: the irreversibility at the abstract, multiplexed operational level ($\Phi$) cannot exceed the total irreversibility of the open physical system ($X$).
2.3.2 Conservation-Congruent Encoding (CCE)
The trajectory-level framework developed in §2.3.1 characterises information in terms of dynamically induced equivalence classes, grouping microscopic configurations that yield statistically indistinguishable macroscopic trajectories under a chosen multiplexed coarse-graining $\Phi$. These equivalence classes are abstract objects; they specify which internal histories are informationally indistinguishable across various macroscopic channels, but do not, by themselves, determine how such distinctions are physically realised. In this subsection, we identify conditions under which these abstract equivalence classes admit metastable physical realizations, making them easier to measure and study. We refer to such realizations as conservation-congruent encodings.
Let $X$ denote the internal physical state space of the agent and let $C$ denote a set of externally adjustable control parameters (e.g., barrier heights, biases, or couplings). The chosen projection $\Phi$ maps microscopic states to a composite quotient space $Z=Z_{1}\times Z_{2}\times\dots\times Z_{C}$, whose elements label the abstract equivalence classes induced by the dynamics across distinct physical channels. A physical system realises a composite coarse-grained state $z=(z_{1},\dots,z_{C})\in Z$ through a probability distribution over its internal physical states under a given control configuration $c\in C$. This is captured by a representation map
$f:Z\times C\to\mathcal{P}(X),$ (24)
where $f(z,c)$ denotes the distribution over microscopic configurations $x\in X$ corresponding to the realization of the joint macroscopic state $z$ under control $c$. At this level, no assumptions are made regarding the stability, uniqueness, or cost of such realizations.
To connect abstract equivalence classes to explicit physical cost, we impose the Conservation-Congruent Encoding (CCE) conditions, which specify when the realizations $f(z,c)$ are metastable and controllably manipulable:
@@ 2.3.2 Conservation-Congruent Encoding (CCE)
    Metastable realization: Each composite equivalence class $z\in Z$ is realised by a metastable basin $B_{z}\subset X$, such that basins corresponding to distinct $z$ are separated by activation barriers associated with the effective, scale-dependent conserved physical quantities that emerge under the chosen projection $\Phi$. Local equilibration within each basin occurs on timescales short compared to those governing inter-basin transitions or control-induced deformation of the basin structure [14, 12].
@@
    Quasistatic, conservation-respecting control: Control trajectories $c_{t}$ deform the basin structure $\{B_{z}\}$ quasistatically relative to internal relaxation. Transitions between basins proceed along paths constrained by the conserved-quantity structure of the dynamics, rather than through arbitrary or nonphysical jumps [7, 21].
Under the CCE assumptions, the representation map $f$ is locally invertible on each metastable basin: physical configurations identify equivalence classes up to fluctuations set by the characteristic scales $\alpha_{c}$ of the specific environmental baths to which the channel couples (e.g., $k_{B}$ for thermal channels, or corresponding intensive parameters for chemical or mechanical channels). As a result, operations that irreversibly merge dynamically distinguishable equivalence classes necessarily entail irreversible physical transformations of these scale-dependent conserved quantities.
It is critical to explicitly distinguish the physical state variables that instantiate an encoding from the conserved quantities exported to erase it. Under CCE, these need not be the same physical property. For example, in a classical CMOS gate, the information is encoded electromagnetically via spatial charge accumulation, but the conserved quantity exported during irreversible erasure is thermal energy (heat). The encoding is termed conservation-congruent because its metastable geometry—the energetic barrier separating its states—is strictly stabilised by, and mathematically coupled to, the specific conservation laws governing the exhaust channel. The irreversible collapse of the encoding state ($x$) strictly forces a proportional export of entropy via the specific conserved quantities via the dissipative flow.
The CCE conditions are sufficient but not necessary for irreversibility to entail entropy export. Crucially, this relationship is scale-dependent. At the absolute microscopic limit of a closed system, dynamic evolution is purely reversible and volume-preserving. Irreversibility and generalised dissipation strictly emerge only at mesoscopic or macroscopic scales where the system is open, requiring the export of conserved quantities into unresolved environmental degrees of freedom. Therefore, for open systems at these higher scales, the collapse of dynamically induced equivalence classes inevitably requires dissipation into degrees of freedom not retained by the system, even when no clean metastable CCE description exists. The CCE framework identifies a broad and practically important class of systems—particularly engineered and experimentally tractable ones—for which this macroscopic cost can be more easily isolated and measured.
In complex systems, information processing is often multiplexed across multiple spatiotemporal scales, meaning distinct conservation-congruent encodings can simultaneously emerge under different coarse-grainings. Consequently, finding strict joint metastability across all physical channels within the raw state space is highly improbable. In such multi-scale regimes, the relevant effective conserved quantities (and their corresponding metastable basins) can be formally understood as distinct slow eigenfunctions of the system’s underlying transfer operator, each associated with a different relaxation timescale. The chosen projection $\Phi$ thus acts as a tunable filter: it sets the specific scale of observation, isolating the functionally relevant metastable encodings at that level from both faster, high-entropy microscopic fluctuations and slower background drift.